# Feynman Integral Reduction without Integration-By-Parts

### v1

@@ 3.4.2 Dimensional recurrence relations
Applying the above relation, we can continue the reduction of $I(-1,2)$ starting from Eq. (79). We first reduce the right-hand side of Eq. (79) to corner integrals in $d+2$ dimensions, and obtain
@@ 4 Examples
[…]
@@
$I(1,2,1,1,1)=4I(3,0,1,1,1)-2I(3,1,0,1,1)-I(2,2,0,1,1)\,.$ (93)
@@ 5 Summary and outlook
Figure 5: The sunrise diagram with $p^{2}=s$.
Given the success in one-loop problems, it is natural to extend our approach to multi-loop integrals. A new ingredient at higher loops is that there can be more than one MIs within a sector, and it is no longer guaranteed that all indices can be reduced to $1$ or 00. Correspondingly, in our approach, we find that when we perform the variable transformations, certain variables appear in the denominators of the transformed integrands. Such transformations are therefore only valid if the corresponding indices are greater than $1$. We have made initial attempts in the sunrise families shown in Figure 5. In the massless case $m_{1}=m_{2}=m_{3}=0$, we can derive reduction relations such as
$I(1,2,2)=\frac{10-3d}{2s}\left[I(1,2,1)+I(1,1,2)\right].$ (94)
Note that in the massless case, $I(1,2,1)$ and $I(1,1,2)$ can be further reduced to $I(1,1,1)$. This is one of the degenerate limits, which have a much richer structure than the one-loop cases. It deserves future investigation how to handle all possible degenerate limits beyond one-loop.
An additional higher-loop feature is that the relations that we can derive by naively extending the one-loop method often involve more than one integrals with the highest complexity. This means that we cannot derive a reduction rule from a single equation, but need to combine several equations. As an example, for the all-massive sunrise family where $m_{1}$, $m_{2}$ and $m_{3}$ are different and larger than zero, we may derive equations such as
$2m_{1}^{2}\left[I(2,2,1)+I(2,1,2)\right]+2m_{2}^{2}\,I(1,3,1)+2m_{3}^{2}\,I(1,1,3)\\ +\left(m_{1}^{2}+m_{2}^{2}+m_{3}^{2}-s\right)I(1,2,2)=\frac{3d-10}{2}\left[I(1,2,1)+I(1,1,2)\right]-I(0,2,2)\,.$ (95)
By solving such a system of equations, we can obtain the desired reduction rules. Note that the number of equations generated in our approach is generically much smaller than that of the usual IBP relations. Furthermore, in our initial attempts we have only naively applied variable transformations similar to the one-loop method. There are still significant flexibility in the integration contours yet to be fully explored. We believe that the system of equations can be further simplified, and it remains possible that one may derive straightforward reduction rules for higher-loop integrals as well.
Finally, in this work we have only exploited the equivalence of integration contours in a simple way. There can be deeper mathematical structures behind these equivalence relations. For example, it is possible to employ the concept of intersection numbers between two integration contours to directly compute the reduction coefficients. This provides a particularly intriguing future perspective.

### v2

@@ 3.4.2 Dimensional recurrence relations
Applying the above relation, we can continue the reduction of I(1,2)I(-1,2) starting from Eq. (81). We first reduce the right-hand side of Eq. (81) to corner integrals in $d+2$ dimensions, and obtain
@@ 4.1 One-loop integrals
$I(1,2,1,1,1)=\left(d-6\right)\frac{\det(\bm{M})}{2\det(\bm{Z})}I(1,1,1,1,1)+\text{sub-sector integrals}\,$, (92)
@@
I(1,2,1,1,1)=4I(3,0,1,1,1)2I(3,1,0,1,1)I(2,2,0,1,1).I(1,2,1,1,1)=4I(3,0,1,1,1)-2I(3,1,0,1,1)-I(2,2,0,1,1)\,. (95)
4.2 A two-loop example
Figure 5: The sunrise diagram with $p^{2}=s$.
Given the success in one-loop problems, it is natural to extend our approach to multi-loop integrals. A new ingredient at higher loops is that there can be more than one MIs within a sector, and it is no longer guaranteed that all indices can be reduced to $1$ or 0. Correspondingly, in our approach, we find that when we perform the variable changes, certain variables may appear in the denominators of the transformed integrands. Such transformations are therefore only valid if the corresponding indices are greater than $1$. In addition, the structure of degenerate limits is also much more complicated than the one-loop case.
As a simple example, consider the massless sunrise family shown in Figure 5, with $m_{1}=m_{2}=m_{3}=0$. The integrals in this family can be represented as
$I(\nu_{1},\nu_{2},\nu_{3})=C(\nu_{1},\nu_{2},\nu_{3})\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{1}\mathrm{d}\alpha_{2}\mathrm{d}\alpha_{3}\,\alpha_{0}^{\nu_{0}-1}\alpha_{1}^{\nu_{1}-1}\alpha_{2}^{\nu_{2}-1}\alpha_{3}^{\nu_{3}-1}\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\left[(\alpha_{1}\alpha_{2}+\alpha_{1}\alpha_{3}+\alpha_{2}\alpha_{3})\alpha_{0}-\alpha_{1}\alpha_{2}\alpha_{3}s\right]^{\lambda_{0}}\,.$ (96)
Following the method in Section 3.2, we define the auxiliary integral
$g(\nu_{1},\nu_{2},\nu_{3})=C(\nu_{1},\nu_{2},\nu_{3})\int_{0}^{\infty}\mathrm{d}\alpha_{0}\mathrm{d}\alpha_{1}\mathrm{d}\alpha_{2}\mathrm{d}\alpha_{3}\,\alpha_{0}^{\nu_{0}-1}\alpha_{1}^{\nu_{1}-1}\alpha_{2}^{\nu_{2}-1}\alpha_{3}^{\nu_{3}-1}\mathcal{X}\delta(\mathcal{X}-\mathcal{Y})\\ \times\left[(\alpha_{1}\alpha_{3}+\alpha_{2}\alpha_{3})\alpha_{0}\right]^{\lambda_{0}}\,$, (97)
which is itself a scaleless integral and should therefore vanish. By performing a variable change $\alpha_{1}\to\alpha_{1}(1+\alpha_{2}/\alpha_{3}-s\alpha_{2}/\alpha_{0})\,$, we can express $g(\nu_{1},\nu_{2},\nu_{3})$ in terms of integrals in the top sector, thus establishing a reduction relation for top-sector integrals. Unlike the case in one-loop integrals, the variable change now introduces the variable $\alpha_{3}$ in the denominator. As a result, the auxiliary integral must have $\nu_{3}\geq 2$ in order to yield a valid reduction relation. For instance, from the auxiliary integral $g(1,1,2)$, we obtain the correct reduction relation
$I(1,2,2)=\frac{10-3d}{2s}\left[I(1,2,1)+I(1,1,2)\right].$ (98)
From IBP reduction, we know that $I(1,2,1)$ and $I(1,1,2)$ can be further reduced to $I(1,1,1)$. Such relations cannot be obtained through the auxiliary integral (97). In fact, this is one of the degenerate limits, which deserves a systematic investigation in the future. Furthermore, in the above example we have only applied variable transformations similar to the one-loop method. There are still significant flexibility in the integration contours yet to be fully explored.
@@ 5 Summary and outlook
We have also demonstrated the validity of our method at two loops. In an explicit two-loop example, the straightforward application of the one-loop method can yield the correct reduction relations. However, we also emphasize that not all reduction relations can be obtained in this way, since additional structures arise at higher loops. This warrants further investigation, including more ways to perform variable changes and contour deformations. Furthermore, in this work we have only exploited the equivalence of integration contours in a simple way. There can be deeper mathematical structures behind these equivalence relations. For example, it is possible to employ the concept of intersection numbers between two integration contours to directly compute the reduction coefficients. This provides a particularly intriguing future perspective.